OpenStudy (anonymous):

find the value of x so that the line passing through (-2 7) and (x 2) has a slope of -1/3

OpenStudy (anonymous):

(-1/3) = (7-2)/(-2-x) Solve for x.

OpenStudy (anonymous):

If you need help just say so and I can walk you through it.

OpenStudy (anonymous):

I could use some help with that..

OpenStudy (anonymous):

Alright well with the equation I provided the first thing we want to do is multiply both sides by (-2-x). So we get something that looks like (-1/3)*(-2-x) = 7-2 Use distributive property on (-1/3)*(-2-x). Do this and then tell me what you get. If you are unable to do this then I will explain it for you.

OpenStudy (anonymous):

Uuh alright well you want me to just answer it for you?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Alright. Well when you distribute it out you should get (2/3) + (x/3) = 7-2 (2/3) + (x/3) = 5 Minus (2/3) from both sides so you get x/3 = 13/3 Multiply both sides by 3 and you get x=39/3 which equals 13. So x = 13. To confirm that this is the correct answer, use the slope formula m = (7-2)/(-2-13) which comes out with m= -1/3 which is the correct answer.

OpenStudy (anonymous):

thank you!